OpenStudy (anonymous):

what is 11/3 simplifyed

OpenStudy (solomonzelman):

\[\huge\color{blue}{ A\frac{b}{c}= \frac{(A \times c) +b}{c} }\] (going the other way,) \[\huge\color{blue}{ \frac{11}{3}= \frac{(3 \times 3) +2}{3}=3\frac{2}{3} }\]

OpenStudy (anonymous):

\[3\frac{ 2 }{ 3 } = \frac{ 11 }{ 3 }\div3 into 11\]